OpenStudy (anonymous):

How many cubic centimeters of a 25% acid solution should be added to 11ccm of a 60% acid solution in order to get a 30% acid solution?

OpenStudy (anonymous):

25(x) + 60(11) = 30(x+11) x =?

OpenStudy (anonymous):

66 THANKS